Q-2): Find the solution of the differential equation: y – 10y + 25y = 2 e5x + sin(x)
Answer #1

SKUHMHNNN YLL loy +254 = 215 + sment AE. D²-10D + 25-01 Dad (P-5)² = O D-55. CF = (G+Goc) est. € P.I = 2 ex-SM (c) D²100 +2